OpenStudy (meilendurcer):

Which of the following is the best starting point for approximating the square root of seventy-three to the nearest whole number?

HanAkoSolo (jamierox4ev3r):

what are the choices for this problem @meilendurcer ?

OpenStudy (anonymous):

73 between what two perfect squares?

OpenStudy (meilendurcer):

\[\sqrt{1}<\sqrt{73}<\sqrt{100}\] \[\sqrt{64}<\sqrt{73}<\sqrt{81}\] \[\sqrt{72}<\sqrt{73}<\sqrt{74}\] \[\sqrt{70}<\sqrt{73}<\sqrt{80}\]

HanAkoSolo (jamierox4ev3r):

the second choice, because we already know the values of the square roots of 64 and of 81

OpenStudy (meilendurcer):

Thank You

HanAkoSolo (jamierox4ev3r):

no problem :)